The population of a town grows at a rate proportional to the

population present at time t. The initial population of 700 increases by 20% in 15 years. What

will be the population in 35 years? How fast is the population growing at t=35?


1
Expert's answer
2021-02-24T07:45:15-0500

The population "P" after "t" years obeys the differential equation

"\frac{dP}{t}=kP;"

where "k" is a positive castant; the initial condition is "P(0)=700".

To solve this, we use separation of variables:

"\int\frac{1}{P}dP=\int kdt"

"ln|P|=kt+C"

"|P|=e^Ce^{kt}"

"P=Ae^{kt}."

Using "P(0)=700" gives "700=Ae^0\Rarr A=700".

Thus, "P=700e^{kt}." Furthermore, "P(15)=700\times120 \%=840" so

"840=700e^{15k}"

"e^{15k}=1.2"

"15k=ln(1.2)"

"k=\frac{ln(1.2)}{15}\approx 0.012."

Thus, "P=700e^{0.012t}." The population after 35 years is therefore

"P=700e^{0.012(35)}=1065.37\approx 1065" people.

Rate of population growth is

"\frac{dP}{t}=\frac{d(700e^{0.012t})}{dt}=(700*0.012)e^{0.012t}"

putting "t=35", we get

"\frac{dP}{t}=(700*0.012)e^{0.012(35)}\approx 13" people/year.